OpenStudy (123ab456c):

A salesperson at a furniture store earns a 3% commission on all sales. How much commission does the salesperson earn on a $1400 sale? A) $21 B) $42 C) $420 D) $1,420

OpenStudy (greatlife44):

you can set up a proportion like this \[\frac{ commission }{ total } = \frac{ percent}{ 100 }\] \[\frac{ x }{ 1400 } = \frac{ 3 }{ 100 }\] then you would cross multiply